OpenStudy (gabylovesyou):

If g(x) = 2(x − 4), find the value of x if g(x) = 20.

hartnn (hartnn):

20=2(x-4) so 10=x-4 can u solve further for x?

OpenStudy (gabylovesyou):

x = 14??

hartnn (hartnn):

yes :)